OpenStudy (anonymous):

Simplify 2/ Radical 5

OpenStudy (anonymous):



OpenStudy (michele_laino):

Please multiply both numerator and denominator by sqrt(5): \[\frac{2}{{\sqrt 5 }} \times \frac{{\sqrt 5 }}{{\sqrt 5 }} = ...?\]

OpenStudy (butterflydreamer):